# RiemannFormer: A Framework for Attention in Curved Spaces

### v2

@@ 2.3.2 $\mathbf{SO}(3)$ Subspaces
where $a^{2}+b^{2}+c^{2}=1$, i.e., $\mathbf{v}=[a,b,c]^{T}$ is a unit vector representing the rotation axis. For a real number $\theta$, the exponential of the antisymmetric matrix $\theta K$ is:
@@ 2.3.3 The Isoclinic Decomposition of $\mathbf{SO}(4)$
We introduce rotation angles $\theta_{A}$ and $\theta_{B}$ for the two types of rotations. The rotation matrices corresponding to positions $m$ and $n$ are then respectively defined as:
@@
$\mathbf{T}_{(x_{1},y_{1})}{\mathbf{T}_{(x_{2},y_{2})}}^{\!-1}$ $=s_{x_{1}}^{-1}s_{y_{1}}^{-1}R(\theta_{A},\theta_{B},(x_{1},y_{1}))s_{x_{2}}s_{y_{2}}R(\theta_{A},\theta_{B},(x_{2},y_{2}))^{-1}$
$={\frac{s_{x_{2}}s_{y_{2}}}{s_{x_{1}}s_{y_{1}}}}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right).$
If we define $\mathbf{M}_{(x_{1},y_{1})}=s^{2}_{x_{1}}s^{2}_{y_{1}}\mathbf{I}$, then the interaction term can be expressed as:
$\mathbf{q}_{\mathbf{m}}^{\top}\mathbf{M}_{\mathbf{m}}\mathbf{T}_{\mathbf{m}}{\mathbf{T}_{\mathbf{n}}}^{\!-1}\mathbf{k}_{\mathbf{n}}$ $=s_{x_{1}}s_{y_{1}}s_{x_{2}}s_{y_{2}}\mathbf{q}_{m}^{\top}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right)\mathbf{k}_{\mathbf{n}}$
@@ A Concrete Realization: The Locality Focusing Mechanism
$\mathbf{o}_{m}=\text{Attention}(\mathbf{q}_{m},\mathbf{K},\mathbf{\Lambda}_{m},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{q}_{m}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{\Lambda}_{m}\mathbf{V})=\mathbf{s}_{m}(\mathbf{\Lambda}_{m}\mathbf{V}),$ (12)
where $\mathbf{s}_{m}\in\mathbb{R}^{1\times L}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times D}$, and $\mathbf{\Lambda}_{m}\in\mathbb{R}^{L\times L}$ is a diagonal matrix containing the factors $\{\lambda_{m,i}\}_{i=1}^{L}$. This expands to:
$\mathbf{o}_{m}$ $=\begin{bmatrix}s_{m,1}&s_{m,2}&\cdots&s_{m,L}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}&&&\\ &\lambda_{m,2}&&\\ &&\ddots&\\ &&&\lambda_{m,L}\end{bmatrix}\begin{bmatrix}\mathbf{v}_{1}\\[3.6pt] \mathbf{v}_{2}\\ \vdots\\ \mathbf{v}_{L}\end{bmatrix}$ (13)
$=\begin{bmatrix}s_{m,1}&s_{m,2}&\cdots&s_{m,L}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}\mathbf{v}_{1}\\[3.6pt] \lambda_{m,2}\mathbf{v}_{2}\\ \vdots\\ \lambda_{m,L}\mathbf{v}_{L}\end{bmatrix}$.
@@
$\mathbf{o}_{m}=\sum_{n=1}^{L}\underbrace{G_{s}(\mathbf{q}_{m},\mathbf{k}_{n})}_{\text{Intensity Similarity}}\underbrace{G_{\sigma}(\mathbf{d}_{m},\mathbf{d}_{n})}_{\text{Spatial Proximity}}\mathbf{v}_{n},$ (14)
@@
For efficient implementation across all tokens, the attenuation factors for all query positions can be collected into a single $L\times L$ matrix $\mathbf{\Omega}$, where $\mathbf{\Omega}_{m,i}=\lambda_{m,i}$. The full attention output is then computed as:
$\mathbf{O}=(\mathbf{S}\odot\mathbf{\Omega})\mathbf{V}=\left(\begin{bmatrix}s_{1,1}&s_{1,2}&\cdots&s_{1,L}\\ s_{2,1}&s_{2,2}&\cdots&s_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ s_{L,1}&s_{L,2}&\cdots&s_{L,L}\end{bmatrix}\odot\begin{bmatrix}\lambda_{1,1}&\lambda_{1,2}&\cdots&\lambda_{1,L}\\ \lambda_{2,1}&\lambda_{2,2}&\cdots&\lambda_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{L,1}&\lambda_{L,2}&\cdots&\lambda_{L,L}\end{bmatrix}\right)\cdot\mathbf{V}.$ (15)
@@ 3 Experiments
In the context of image data, there are two axes to consider. Although these axes can be addressed jointly [heo2024ropevit] or the image patches are directly flatten into a sequence, we are interested in exploring the effects of processing them separately along the $x$ and $y$ axes. This strategy also facilitates a comparative studies with previous work.
The proposed modules undergo training and evaluation on the CIFAR-10 and CIFAR-100 datasets, employing a lightweight Vision Transformer (ViT) architecture. Specifically, these modules leverage the compact yet efficient structure of the ViT model to process the image classification tasks inherent to these benchmark datasets.
All experiments were implemented on an NVIDIA RTX 4090 GPU with PyTorch 2.7. The training procedure is conducted over 200 epochs with a batch size of 128, utilizing the AdamW optimizer configured with its default hyperparameters.
In predefined transformations, rotation operations in subspaces involve the configuration of angles. An angle schedule is a function that sends an integer to an angle involved in the linear transformation. The choice of the angle schedule can also impact the performance of the tangent mapping. In our experiments, we observed that the following straightforward schedule yields favorable results: $f(t):t\mapsto t\theta,\;t\in\{1,2,\dots\}$, where $\theta$ is a learnable parameter or a predefined angle. Exploring more effective angle schedules is also a direction worth investigating. In the current experiment of this work, we use 1D PaTE-SO(2) with the Conformal Exponential Scaling scheme ($\mathbf{M}_{i}=s^{i}\mathbf{I}$), and we set the hyperparameter $\alpha=0.1$ in $s=\frac{\exp(w)}{\exp(w)+\alpha}$ in these experiments.
@@
Table 1: Comparison with absolute/relative positional encodings on small datasets in image classification benchmarks.
Datasets No-Pos Sinusoidal RoPE LF only RiemannFormer RiemannFormer+LF
CIFAR-10 79.61 86.06 87.91 87.30 89.87 90.82
CIFAR-100 53.37 59.81 62.37 63.01 64.51 66.96
In Table 1, LF is an abbreviation for Locality Focusing mechanism. When employing solely LF without any positional encoding, the module exhibits mediocre performance; however, incorporating LF into positional encoding results in a significant enhancement in model performance. This suggests that LF plays a crucial role in improving performance on these small datasets, whereas positional encoding alone lacks the inherent properties of LF.
@@ 4 Conclusion
We assume that the data resides in a curved space, and allow the model to adaptively explore the geometric structure of the space from the data itself, thereby obviating the need for explicit positional information. Our framework can capture both absolute and relative position information and allows the model autonomously balance them during the learning process. Through predefined configurations, we guide the model to capture the geometric structure of curved space while simultaneously reducing the number of parameters that need to be learned. In addition, we introduce a decay factor to attenuate the influence of remote positions. This mechanism facilitates further our modules achieve substantial performance improvements over the baseline.
The framework outlined in this paper has been demonstrated to effectively improve the performance of transformers by conducting experiments on the CIFAR-10 and CIFAR-100 datasets. The current implementation in this paper serves as a preliminary validation of our insights, and further evaluation experiments involving more visual models and large language models will be conducted sequentially.
Appendix
We can also learn a metric matrix of general form, which is not necessarily restricted to the diagonal matrix. For clarity and simplicity, we present here only the $2\times 2$ matrix corresponding to two-dimensional subspace.
Let,
$\mathbf{S}=\left[\begin{matrix}s_{1}&0\\[8.09998pt] 0&s_{2}\end{matrix}\right],$
where $s_{1}>0$ and $s_{2}>0$. Then, the symmetric positive definite matrices are preset as follows,
$\mathbf{M}_{m}$ $=\left[\begin{matrix}\cos{\left(m\theta\right)}&-\sin{\left(m\theta\right)}\\[8.09998pt] \sin{\left(m\theta\right)}&\;\;\;\cos{\left(m\theta\right)}\end{matrix}\right]^{\top}\left[\begin{matrix}s_{1}^{m}&0\\[8.09998pt] 0&s_{2}^{m}\end{matrix}\right]\left[\begin{matrix}\cos{\left(m\theta\right)}&-\sin{\left(m\theta\right)}\\[8.09998pt] \sin{\left(m\theta\right)}&\;\;\;\cos{\left(m\theta\right)}\end{matrix}\right]$
$=\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{m}\cdot\breve{\bf R}(m\theta),$
and,
$\mathbf{M}_{n}$ $=\left[\begin{matrix}\cos{\left(n\theta\right)}&-\sin{\left(n\theta\right)}\\[8.09998pt] \sin{\left(n\theta\right)}&\;\;\;\cos{\left(n\theta\right)}\end{matrix}\right]^{\top}\left[\begin{matrix}s_{1}^{n}&0\\[8.09998pt] 0&s_{2}^{n}\end{matrix}\right]\left[\begin{matrix}\cos{\left(n\theta\right)}&-\sin{\left(n\theta\right)}\\[8.09998pt] \sin{\left(n\theta\right)}&\;\;\;\cos{\left(n\theta\right)}\end{matrix}\right]$
$=\breve{\bf R}(-n\theta)\cdot\mathbf{S}^{n}\cdot\breve{\bf R}(n\theta).$
Then, the linear transformations can be preset as the following matrices,
$\mathbf{T}_{m}$ $=\left[\begin{matrix}\;\;\;\cos{\left(m\theta\right)}&\sin{\left(m\theta\right)}\\[8.09998pt] -\sin{\left(m\theta\right)}&\cos{\left(m\theta\right)}\end{matrix}\right]\left[\begin{matrix}s_{1}^{-\frac{m}{2}}&0\\ 0&s_{2}^{-\frac{m}{2}}\end{matrix}\right]$
$=\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{-\frac{m}{2}},$
and,
$\mathbf{T}_{n}$ $=\left[\begin{matrix}\;\;\;\cos{\left(n\theta\right)}&\sin{\left(n\theta\right)}\\[8.09998pt] -\sin{\left(n\theta\right)}&\cos{\left(n\theta\right)}\end{matrix}\right]\left[\begin{matrix}s_{1}^{-\frac{n}{2}}&0\\ 0&s_{2}^{-\frac{n}{2}}\end{matrix}\right]$
$=\breve{\bf R}(-n\theta)\cdot\mathbf{S}^{-\frac{n}{2}}.$
It is straightforward to verify that:
$\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}=\mathbf{T}_{n}^{\top}\mathbf{M}_{n}\mathbf{T}_{n}.$
and,
$\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$ $=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{m}\cdot\breve{\bf R}(m\theta)\cdot\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{-\frac{m}{2}}\cdot\mathbf{S}^{\frac{n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{\frac{m+n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}.$
Then, the form of a standard inner product is,
$\langle\mathbf{S}^{\frac{m}{2}}\breve{\bf R}(m\theta)\mathbf{q}_{m},\;\mathbf{S}^{\frac{n}{2}}\breve{\bf R}(n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$

### v3

@@ 2.4.2 $\mathbf{SO}(3)$ Subgroups
where $a^{2}+b^{2}+c^{2}=1$, i.e., $\mathbf{v}=[a,b,c]^{\top}$ is a unit vector representing the rotation axis. For a real number $\theta$, the exponential of the antisymmetric matrix $\theta K$ is:
@@ 2.4.3 The Isoclinic Decomposition of $\mathbf{SO}(4)$
Extension to Block-Diagonal Rotations in $\mathbf{SO}(4k)$
The powerful decomposition of $\mathfrak{so}(4)$ provides a clear blueprint for constructing and parameterizing a significant subgroup of higher-dimensional rotations within $\mathbf{SO}(4k)$. We can generalize this framework by considering the Lie algebra of block-diagonal skew-symmetric matrices, which generates a corresponding block-diagonal subgroup of $\mathbf{SO}(4k)$.
Let us define a specific Lie subalgebra of $\mathfrak{so}(4k)$ as the direct sum of $k$ copies of $\mathfrak{so}(4)$:
$\mathfrak{g}=\bigoplus_{i=1}^{k}\mathfrak{so}(4)^{(i)}\subset\mathfrak{so}(4k).$
An element $H\in\mathfrak{g}$ is a $4k\times 4k$ block-diagonal matrix, where each block is an element of $\mathfrak{so}(4)$:
$H=\operatorname{diag}(X_{1},X_{2}\cdots,X_{k}),$
with each $X_{i}\in\mathfrak{so}(4)$. The off-diagonal blocks are all zero matrices.
A fundamental property of the matrix exponential is that it preserves this block-diagonal structure. The exponential of $H$ is simply the block-diagonal matrix of the individual matrix exponentials:
$\exp(H)=\operatorname{diag}(\exp(X_{1}),\exp(X_{2}),\cdots,\exp(X_{k}))$
Since the exponential map takes an element from the Lie algebra $\mathfrak{so}(4)$ to the Lie group $\mathbf{SO}(4)$, each block $\exp(X_{i})$ is a rotation matrix in $\mathbf{SO}(4)$. Consequently, $\exp(H)$ is an element of $\mathbf{SO}(4k)$, and the set of all such matrices forms a Lie subgroup that is the direct product of $k$ copies of $\mathbf{SO}(4)$, embedded within $\mathbf{SO}(4k)$.
We can now apply the isoclinic decomposition to each block $X_{i}$ independently. For each $i\in\{1,\dots,k\}$, we decompose $X_{i}$ into its left-isoclinic and right-isoclinic components:
$X_{i}=A_{i}+B_{i},$
where $A_{i}\in\mathfrak{h}_{1}^{(i)}$ and $B_{i}\in\mathfrak{h}_{2}^{(i)}$. Here, $\mathfrak{h}_{1}^{(i)}$ and $\mathfrak{h}_{2}^{(i)}$ are the commuting subalgebras of the $i$-th copy of $\mathfrak{so}(4)$.
Substituting this back into the block-diagonal matrix $H$, we can decompose $H$ into two distinct block-diagonal matrices:
$H=H_{A}+H_{B},$
where
$H_{A}=\operatorname{diag}(A_{1},A_{2},\dots,A_{k})\quad\text{and}\quad H_{B}=\operatorname{diag}(B_{1},B_{2},\dots,B_{k}).$
The matrix $H_{A}$ aggregates all the left-isoclinic components, while $H_{B}$ aggregates all the right-isoclinic components.
Crucially, these two matrices commute, $H_{A},H_{B}]=0$. This is because the commutator of block-diagonal matrices is the block-diagonal matrix of the commutators of their corresponding blocks:
$H_{A},H_{B}]=\operatorname{diag}([A_{1},B_{1}],[A_{2},B_{2}],\dots,[A_{k},B_{k}]).$
As established earlier, $A_{i},B_{i}]=0$ for all $i$, so every block in the resulting matrix is a zero matrix.
This commutativity allows for the final factorization of the exponential map for the entire block-diagonal rotation in $\mathbf{SO}(4k)$:
$R=\exp(H)=\exp(H_{A}+H_{B})=\exp(H_{A})\exp(H_{B}).$
The resulting rotation $R$ is the product of two commuting block-diagonal rotations, $R_{A}=\exp(H_{A})$ and $R_{B}=\exp(H_{B})$. Each of these can be computed efficiently block by block using the closed-form Euler-like formula derived previously:
$R_{A}=\operatorname{diag}(\exp(A_{1}),\dots,\exp(A_{k}))=\operatorname{diag}(\cos(\theta_{1,i})I+\sin(\theta_{1,i})J_{A_{i}})_{i=1}^{k},$
$R_{B}=\operatorname{diag}(\exp(B_{1}),\dots,\exp(B_{k}))=\operatorname{diag}(\cos(\theta_{2,i})I+\sin(\theta_{2,i})J_{B_{i}})_{i=1}^{k}.$
This framework effectively reduces the complex problem of parameterizing a $4k\times 4k$ rotation within this subgroup to the management of $2k$ independent angles and $2k$ independent 3D axes, offering a computationally tractable and structurally elegant solution.
To apply this machinery in a practical context, such as modeling sequential data where each block corresponds to a specific position, we can introduce a more structured parameterization. Instead of learning independent angles for each position, we can define the rotation for a given position as a function of its index. This enforces a consistent geometric relationship across the sequence.
We introduce global base rotation angles $\theta_{A}$ and $\theta_{B}$ for the two types of isoclinic rotations. The rotation matrices corresponding to positions $m$ and $n$ are then respectively defined as:
@@
$\mathbf{T}_{(x_{1},y_{1})}{\mathbf{T}_{(x_{2},y_{2})}}^{\!-1}$ $={f(x_{1},y_{1})}^{-1}R(\theta_{A},\theta_{B},(x_{1},y_{1}))f(x_{2},y_{2})R(\theta_{A},\theta_{B},(x_{2},y_{2}))^{-1}$
$={\frac{f(x_{2},y_{2})}{f(x_{1},y_{1})}}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right).$
If we define $\mathbf{M}_{(x_{1},y_{1})}=f(x_{1},y_{1})^{2}\mathbf{I}$, then the interaction term can be expressed as:
$\mathbf{q}_{\mathbf{m}}^{\top}\mathbf{M}_{\mathbf{m}}\mathbf{T}_{\mathbf{m}}{\mathbf{T}_{\mathbf{n}}}^{\!-1}\mathbf{k}_{\mathbf{n}}$ $=f(x_{1},y_{1})f(x_{2},y_{2})\mathbf{q}_{m}^{\top}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right)\mathbf{k}_{\mathbf{n}}$
@@ A Concrete Realization: The Locality Focusing Mechanism
$\mathbf{o}_{m}=\text{Attn}(\mathbf{q}_{m},\mathbf{K},\mathbf{\Lambda}_{m},\mathbf{V})=\mathbf{C}_{m}(\mathbf{\Lambda}_{m}\mathbf{V}),$ (14)
where $\mathbf{C}_{m}\in\mathbb{R}^{1\times N}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times D}$, and $\mathbf{\Lambda}_{m}\in\mathbb{R}^{N\times N}$ is a diagonal matrix containing the factors $\{\lambda_{m,i}\}_{i=1}^{N}$. This expands to:
$\mathbf{o}_{m}$ $=\begin{bmatrix}C_{m,1}&C_{m,2}&\cdots&C_{m,N}\end{bmatrix}\operatorname{diag}(\lambda_{m,1},\lambda_{m,2},\cdots,\lambda_{m,N})\begin{bmatrix}\mathbf{v}_{1}\\[3.6pt] \mathbf{v}_{2}\\ \vdots\\ \mathbf{v}_{N}\end{bmatrix}$ (15)
$=\begin{bmatrix}C_{m,1}&C_{m,2}&\cdots&C_{m,N}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}\mathbf{v}_{1}\\[3.6pt] \lambda_{m,2}\mathbf{v}_{2}\\ \vdots\\ \lambda_{m,N}\mathbf{v}_{N}\end{bmatrix}$.
@@
$\mathbf{o}_{m}=\sum_{n=1}^{N}\underbrace{G_{s}(\mathbf{q}_{m},\mathbf{k}_{n})}_{\text{Intensity Similarity}}\underbrace{G_{\sigma}(\mathbf{d}_{m},\mathbf{d}_{n})}_{\text{Spatial Proximity}}\mathbf{v}_{n},$ (16)
@@
For efficient implementation across all tokens, the attenuation factors for all query positions can be collected into a single $N\times N$ matrix $\mathbf{\Omega}$, where $\mathbf{\Omega}_{m,i}=\lambda_{m,i}$. The full attention output is then computed as:
$\mathbf{O}=(\mathbf{C}\odot\mathbf{\Omega})\mathbf{V}=\left(\begin{bmatrix}C_{1,1}&C_{1,2}&\cdots&C_{1,N}\\ C_{2,1}&C_{2,2}&\cdots&C_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ C_{N,1}&C_{N,2}&\cdots&C_{N,N}\end{bmatrix}\odot\begin{bmatrix}\lambda_{1,1}&\lambda_{1,2}&\cdots&\lambda_{1,N}\\ \lambda_{2,1}&\lambda_{2,2}&\cdots&\lambda_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{N,1}&\lambda_{N,2}&\cdots&\lambda_{N,N}\end{bmatrix}\right)\cdot\mathbf{V}.$ (17)
@@ 3 Experiments
Image data inherently possesses a 2D structure defined by two spatial axes. While standard approaches often flatten patches into a 1D sequence or address these dimensions jointly [heo2024ropevit], our PaTE-SO(4) framework offers a mathematically natural correspondence. The algebraic decomposition of SO(4) yields two commuting sets of rotational generators (left- and right-isoclinic rotations). This dual-axis structure aligns perfectly with the Cartesian geometry of images, allowing us to explicitly and simultaneously encode the height and width dimensions within the same unified representation.
We define the Riemannian metric at a patch position $(x,y)$ as $\mathbf{M}_{(x,y)}=f(x,y)^{2}\mathbf{I}$. Adopting PaTE-SO(4) as a representative instantiation, the interaction between two tokens located at different positions can be expressed as:
$\langle\mathbf{T}_{(x_{1},y_{1})}^{-1}\mathbf{q}_{(x_{1},y_{1})},\!\mathbf{T}_{(x_{2},y_{2})}^{-1}\mathbf{k}_{(x_{2},y_{2})}\rangle=f(x_{1},y_{1})f(x_{2},y_{2})\mathbf{q}_{(x_{1},y_{1})}^{\top}\exp\left((x_{1}\!-\!x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}\!-\!y_{2})\theta_{B}J_{B}\right)\mathbf{k}_{(x_{2},y_{2})}$
where $\mathbf{T}_{(x,y)}=f(x,y)^{-1}R(\theta_{A},\theta_{B},(x,y))$ encapsulates both rotation and scaling. Then we decouple the scaling component from the Softmax to form our attention module $\operatorname{RieAttn}=\left(\mathbf{L}\mathbf{L}^{\top}\odot\operatorname{Softmax}\left(\frac{\mathbf{R}[\mathbf{Q}]\,\mathbf{R}[\mathbf{K}]^{\top}}{\sqrt{d_{k}}}\right)\right)\mathbf{V}$.
To empirically validate the efficacy of this geometrically principled mechanism, particularly in capturing intrinsic spatial structures, we integrate it into a practical vision backbone. The proposed modules undergo training and evaluation on the CIFAR-10 and CIFAR-100 datasets, employing a lightweight Vision Transformer (ViT) architecture. Specifically, these modules leverage the compact yet efficient structure of the ViT model to process the image classification tasks inherent to these benchmark datasets. All experiments were implemented on an NVIDIA RTX 4090 GPU with PyTorch 2.7. The training procedure is conducted over 100 epochs with a batch size of 128, utilizing the AdamW optimizer configured with its default hyperparameters.
In our experimental evaluation, we implement and compare both PaTE-SO(2) and PaTE-SO(4) variants. For both formulations, the rotation operations rely on a specific angle configuration across different subspaces. We define an angle schedule function that maps a subspace index to its corresponding rotation angle. Empirically, we observed that a straightforward harmonic schedule yields favorable results: $f(t):t\mapsto t\theta,\;t\in\{1,2,\dots,K\}$, where $\theta$ is a learnable base frequency and $t$ denotes the subspace index.
Furthermore, for the PaTE-SO(4) instantiation, the geometric configuration extends beyond mere angles to the orientation of the rotation axes (specifically, the axes for the left and right isoclinic rotations). To investigate the trade-off between geometric flexibility and parameter efficiency, we evaluate two distinct axis strategies: independent axes, where each subspace learns its own unique rotation axis, and shared axes, where a single global axis orientation is applied across all subspaces.
@@
Table 1: Results of the Proposed Models on Small Datasets in Image Classification Benchmarks.
Datasets Models PaTESO(2) PaTESO(4)
Rot Rot+Scl Rot+Scl+LF Rot Rot+Scl Rot+Scl(SA) Rot+Scl+LF Rot+Scl+LF(SA)
CIFAR-10 82.11 83.68 85.71 84.75 85.44 85.32 85.61 86.10
CIFAR-100 53.49 53.81 57.49 55.43 55.39 57.68 58.75 58.98
Table 1 presents the preliminary experimental results of our proposed algorithms on the CIFAR-10 and CIFAR-100 datasets. For clarity, the abbreviations used in the table are defined as follows: Rot denotes the rotational component; Scl represents the Riemannian metric scaling; LF refers to the Locality Focusing Mechanism; and SA indicates the Shared Axis strategy within the PaTESO(4) framework, where the rotation axes are shared across all subspaces.
A comparative analysis of the tabulated data reveals several key insights into the proposed framework. First, PaTESO(4) consistently outperforms PaTESO(2) across both datasets and configurations (e.g., increasing from 82.11% to 84.75% in the base ’Rot’ setting on CIFAR-10), suggesting that the higher-dimensional rotational structure of SO(4) captures spatial relationships more effectively than SO(2). Second, the integration of Riemannian metric scaling (Scl) yields observable gains over the pure rotation baseline in most cases, validating the hypothesis that modulating the geometry based on content or position enhances representation. Third, the Locality Focusing (LF) mechanism provides a significant performance boost (e.g., reaching 85.71% in PaTESO(2)), highlighting the importance of inductive biases for local neighborhood attention in visual tasks. Finally, the combination of all components yields the highest accuracy, with PaTESO(4) using the Shared Axis (SA) strategy achieving the best overall results (86.10% on CIFAR-10 and 58.98% on CIFAR-100).
It is important to note that the primary objective of these experiments is to evaluate the relative efficacy of different geometric schemes rather than to chase state-of-the-art absolute performance. Consequently, we employed a relatively small network architecture and a limited number of training iterations. For the metric scaling (Scl), we adopted a dynamic approach where the scale is determined by the embeddings using the simple projection scheme defined in Eq. (10). Regarding PaTESO(4), we observed that the Shared Axis (SA) variant appears to perform slightly better than the independent axis configuration under the current angle strategy. We hypothesize that this may be due to the regularization effect of sharing parameters, which is particularly beneficial given the current model scale and parameter constraints. Further investigations on larger datasets and deeper networks are required to fully validate the scalability of the independent axis strategy.
@@ 4 Conclusion
We present RiemannFormer, a framework that reimagines the Transformer through the lens of differential geometry. We assume that the tokens reside in a curved space, and their latent representations are governed by parallel transport and Riemannian metrics. Our approach utilizes predefined Lie group structures to capture spatial dependencies efficiently, removing the reliance on traditional positional embeddings. A key contribution is the Riemannian Attention Mechanism, which strategically decouples metric scaling from softmax normalization. This design allows us to implement a dynamic scaling strategy where the embedding content itself dictates the local geometry — akin to mass curving spacetime. This mechanism not only serves as an effective locality bias by attenuating remote interactions but also provides a content-aware gating capability.
The framework outlined in this paper has been demonstrated to effectively improve the performance of transformers by conducting experiments on the CIFAR-10 and CIFAR-100 datasets. Out experimental results confirm that properly modeling this intrinsic geometry yields significant performance gains, particularly in visual tasks. The current implementation in this paper serves as a preliminary validation of our insights, and further evaluation experiments involving more visual models and large language models will be conducted sequentially.